HUBBLE SPACE TELESCOPE DAILY REPORT #3088 PERIOD COVERED: 0000Z (UTC) 04/04/02 - 0000Z (UTC) 04/05/02 Daily Status Report as of 095/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of ACS/WFC/HRC 8947 (Weekly Test) The Advanced Camera for Surveys (WFC and HRC) was used to perform basic tests to monitor, the read noise, the development of hot pixels and test for any source of noise in ACS CCD detectors. This program will be executed at least once a day for the entire lifetime of ACS. There were no problems reported. 1.2 Completed WF/PC-2 9060 (Photometry of a Statistically Significant Sample of Kuiper Belt Objects) The WF/PC-2 was used to propel the physical study of KBOs forward by performing accurate photometry at V, R, and I on a sample of up to 150 KBOs. The sample is made up of objects that will be observed at thermal infrared wavelengths by SIRTF and will be used with those data to derive the first accurate diameters and albedos for a large sample of KBOs. The observations completed nominally. 1.3 Completed Two Sets of STIS/CCD 9088 (Next Generation Spectral Library of Stars) The Space Telescope Imaging Spectrograph (CCD) was used to produce a "Next Generation'' Spectral Library of 600 stars for use in modeling the integrated light of galaxies and clusters by using the low dispersion UV and optical gratings of STIS. The library will be roughly equally divided among four metallicities, very low {Fe/H < -1.5}, low {-1.5 < Fe/H < -0.5}, near-solar {-0.5 < Fe/H < 0.1}, and super-solar {Fe/H > 0.1}, well-sampling the entire HR-diagram in each bin. Such a library will surpass all extant compilations and have lasting archival value, well into the Next Generation Space Telescope era. No problems occurred. 1.4 Completed Two Sets of WF/PC-2 8937 (Cycle 9 Supplemental Darks pt2/3) The WF/PC-2 was used obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. No problems were encountered. 1.5 Completed STIS/CCD 8904 (Bias Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4 in order to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no anomalous activity. 1.6 Completed FGS/1 9237 (The Masses of the O-type Binary 15 Monocerotis) Fine Guidance Sensor #1 was used to observe the O-type star 15 Mon {HD 47839} that was recently discovered to be an astrometric and spectroscopic binary with an orbital period of 25 years. It is the first known O-star system to bridge the observational gap between the period regimes normally probed by these techniques. The observations completed with no reported problems 1.7 Completed STIS/CCD 9147 (Probing The Structure Of The Shocks In The Narrow Line Region Of M 51) The Space Telescope Imaging Spectrograph (CCD) was used to determine whether nuclear or shock-induced photoionization dominates the gaseous ionization in the narrow-line region {NLR} of LINERs and Seyfert galaxies. A critical issue is the determination of the input of mechanical energy in the NLR through jet/ISM interactions. Previous studies suggest that the latter could actually suffice to power the NLR in most Seyfert galaxies. Previous approaches to this problem have obtained integrated UV/visible line flux ratios, which are sensitive to uncertain reddening. It is now proposed to map several spectral lines across the shock/precursor complex in the spatially extended, jet-driven NLR of the nearby Seyfert 2 galaxy M 51, measuring the spatial distribution of kinematic subsystems, the ionization structure, and the gaseous density, across this structure. All observations completed nominally. 1.8 Completed FGS/1 9034 (The Masses and Luminosities of Population II Stars) Fine Guidance Sensor #1 was used to observe the mass-luminosity relation {MLR} of Population II stars of which very little is currently known. With the advent of the Hipparcos Catalogue, improved distances to many spectroscopic binaries known to be Pop II systems are now available. After surveying the literature and making reasonable estimates of the secondary masses, we find 13 systems whose minimum separation should be larger than the resolution limit of FGS #1. The observations completed nominally. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 7 Successful: 7 Scheduled Re-acquisitions: 10 Successful: 10 2.2 FHST Updates: Scheduled: 16 Successful: 16 2.3 Operations Notes: The NICMOS Cryo Cooler (NCS) cool down continues nominally. As of yesterday morning, the NICMOS Outlet Temperature (MNPNCOLT) was 104K (expected end point 76K), the NCC TA Inlet Temperature (MNPTAILT) was 83K (expected end point 68K), the CLI Temperature (MNPLINTT) was 88K (expected end point 71K), and the NICMOS Inlet Temperature (MNRNCILT) was 78K. Per operations requests, the NCS CPL minimum value was set to 12 at 094/1550Z, the min/max collection for the NCS post-surge period was re-configured at 094/1937Z, and the 8051 flight software end-of-surge value was set to 82 at 094/1938Z. The STIS MCE-1 reset at 094/204158Z while the low voltage was on and while outside any SAA interval. MAMA-1 will be recovered via SMS commanding when the high voltage is turned on. Per ROP NS-12, the STIS flight software error counter was cleared at 094/204158Z. An operations note was utilized to change the MAMA threshold voltage limits. A TTR was written when there was a required re-transmit at 095/0916Z during an NSSC-1 load. SI C&DH errors were reset at 095/0922Z per ROP NS-5. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of Servicing Mission Orbital Verification and the gradual resumption of normal science observations and calibrations. /CAW